OpenStudy (anonymous):

Find the derivative of f(x) = 5/x at x = -1.

OpenStudy (anonymous):

by hand, or can you use the formula for the derivative of \(\frac{5}{x}\) ?

OpenStudy (anonymous):

It doesnt matter

OpenStudy (anonymous):

if you can use the formula it is easy \[\frac{d}{dx}[\frac{1}{x}]=-\frac{1}{x^2}\] and so \[\frac{d}{dx}[\frac{5}{x}]=-\frac{5}{x^2}\]

OpenStudy (anonymous):

put \(x=-1\) and you get the answer

OpenStudy (anonymous):

if you have to do it taking limits, it is a bit harder with the algebra

OpenStudy (anonymous):

Oh, Thanks!

OpenStudy (anonymous):

yw